In each of the following systems , determine the solution ( Find the intersection of both lines
y=2x+3
y=-x+5

Answer by drk(1908)

Using substitution, we get
2x + 3 = -x + 5
subtract x to get
3x + 3 = 5
subtract 3 to get
3x = 2
divide to get
x = 2/3
If x = 2/3 then y = 4 1/3.
Your coordinate of intersection is
(2/3, 4 1/3)